A shopkeeper buys 80 articles for ₹ 2,400 and sells them for a profit of 16%. Find the selling price of one article. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given that the shopkeeper buys 80 articles for₹2,400.
Cost of one article =2400/80 =₹30
Profit percent =Profit× 100
C.P.
16=Profit/ ₹30× 100
Profit =16 ×30
100
=4.8
∴Selling price of one article = C.P. + Profit =₹(30+4.80)
=₹34.80

Thanks